Hi,
Can someone please show me step by step on how to solve this problem?


3(4y+5)=25

Thank You!!

Answer by mananth(16946)

3(4y+5)=25
...
3*4y+3*5=25
12y+15=25
add -15 to both sides
12y+15-15=25-15
12y=10
divide both sides by 12
12y/12 = 10/12
y = 5/6
..
m.ananth@hotmail.ca